OpenStudy (emendez10):

Prove that every tangent plane to the surface z= xe^(x/y) passes through the origin.

OpenStudy (eliesaab):

The equation of the tangent plan at any point (a,b) is given by z= f(a,b) +Gradient(f,(a,b)) dot (x-a,y-b)

OpenStudy (eliesaab):

First you find the gradient vector at (a,b), you find \[ \left\{\frac{a e^{a/b}}{b}+e^{a/b},-\frac{a^2 e^{a/b}}{b^2}\right\} \]

OpenStudy (eliesaab):

Apply the above formula to find \[ z=-\frac{a^2 e^{a/b} (y-b)}{b^2}+\left(\frac{a e^{a/b}}{b}+e^{a/b}\right) (x-a)+a e^{a/b} \]

OpenStudy (eliesaab):

If you put x=0 and y=0 above, you find that z=0, meaning that the the plan passes through the origin

OpenStudy (irishboy123):

you sure? the normal vector to that surface is \(\nabla z = \vec n = <e^{\frac{x}{y}}(1+ \frac{x}{y}), - e^{\frac{x}{y}} (\frac{x}{y})^2>\) \(= e^{\frac{x}{y}} <(1+ \frac{x}{y}), - (\frac{x}{y})^2>\) the equation of any plane is \((\vec r - \vec r_o) \bullet \vec n = 0\) if the origin is on every tangent plane, it follows that for every plane: \((\vec r - \vec 0) \bullet \vec n = 0\) or \( \vec r \bullet \vec n = 0\) yet \(<x, y> \bullet \vec n \) \( = <x, y> \bullet e^{\frac{x}{y}} <(1+ \frac{x}{y}), - (\frac{x}{y})^2> \ne 0\)

OpenStudy (loser66):

@518nad

OpenStudy (irishboy123):

what is z(0,0)

OpenStudy (518nad):

oh u can think about it like this actually we have a parametric equation like this R(x,y)=<x,y,f(x,y)> R_x cross R_y = normal vector at (x,y,z) = <a,b,c> plane equation ax+by+cz=D z=f(x,y) use the 2 equatiosn to get expression for D after subbing this stuff in, you will be left with a plane equation just as a function of x,y,z and if x,y,z=0 = 0 is true then you are done

OpenStudy (irishboy123):

LOL!!!

OpenStudy (518nad):

hey u stop laughing at me

OpenStudy (brooke..help00):

If have answer already please close:) Thanks!

OpenStudy (irishboy123):

Just laughing at meself, mate! I didn't write the closed surface first :-( So one more time: \(S = x e^{x/y} - z = 0\) the normal vector to that surface is \(\vec n = \nabla S = <e^{\frac{x}{y}}(1+ \frac{x}{y}), - e^{\frac{x}{y}} (\frac{x}{y})^2, -1>\) the equation of any plane is \((\vec r - \vec r_o) \bullet \vec n = 0\) if the origin is on every tangent plane, it follows that for every plane: \((\vec r - \vec 0) \bullet \vec n = 0\) or \( \vec r \bullet \vec n = 0\) And \(<x, y, z> \bullet \vec n \) \( = <x, y, z> \bullet <e^{\frac{x}{y}}(1+ \frac{x}{y}), - e^{\frac{x}{y}} (\frac{x}{y})^2, -1> \color{red}{ = 0}\)